If 1 is a zero of p(x) =ax^3-(a-1)x-1 then find value of a
Answer by mananth(16946)


p(x) =ax^3-(a-1)x-1
since 1 is the root
p(1) = a(1)^3-(1-1)x -1
=a-1
by remainder theorem
a-1 = 0
therefore a=1